OpenStudy (anonymous):

Suppose a parabola has vertex (5, –3) and also passes through the point (6, 1). Write the equation of the parabola in vertex form. y = (x – 5)2 – 3 y = 4(x – 5)2 – 3 y = 4(x + 5)2 – 3 y = 4(x – 5)2 + 3

OpenStudy (anonymous):

ok if the vertex is \((5,-3)\) then it looks like \[y=a(x-5)^2-3\]

OpenStudy (anonymous):

then to find out which one it is, you know if \(x=6\) then \(y=1\) because that is what "passes through the point \((6,1)\) means replace \(x\) by \(6\) and set the result equal to \(1\)

OpenStudy (anonymous):

you get \[1=a(6-5)^2-3\] \[1=a-3\] \[4=a\]

OpenStudy (anonymous):

pick \[y = 4(x – 5)^2 – 3\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can you help with more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

identify the vertex and the axis of symmetry of the graph of the function y = 3(x + 2)2 – 3. (1 point) vertex: (2, –3); axis of symmetry: x = 2 vertex: (–2, –3); axis of symmetry: x = –2 vertex: (2, 3); axis of symmetry: x = 2 vertex: (–2, 3); axis of symmetry: x = –2

OpenStudy (anonymous):

\[y=3(x+2)^2-3\] read off from \(y=a(x-h)^2+k\) with vertex \((h,k)\) and axis of symmetry \(x=h\)

OpenStudy (anonymous):

Identify the maximum or minimum value and the domain and range of the graph of the function y = 2(x – 3)2 – 4. (1 point) minimum value: –4 domain: all real numbers range: all real numbers –4 maximum value: 4 domain: all real numbers range: all real numbers 4 maximum value: –4 domain: all real numbers –4 range: all real numbers minimum value: 4 domain: all real numbers4 range: all real numbers

OpenStudy (anonymous):

in your case \(h=-2\) and \(k=-3\) so vertex is \((-2,-3)\) and axis of symmetry is \(x=-2\)

OpenStudy (anonymous):

that was for the second question you posted

OpenStudy (anonymous):

third question Identify the maximum or minimum value and the domain and range of the graph of the function \(y = 2(x – 3)^2 – 4\)

OpenStudy (anonymous):

this is a parabola that opens up, so it has a minimum, not a maximum. the minimum is the second coordinate of the vertex, which is \(-4\) so the range is from \(-4\) to \(\infty\) or "all real numbers greater than or equal to \(-4\)

OpenStudy (anonymous):

\(-4\) is the minimum the domain of any polynomial is all real numbers

OpenStudy (anonymous):

i don't really understand your choices, but if the first one is minimum value: –4 domain: all real numbers range: all real numbers \(\geq –4\) then it is the right one

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The graph below models the path of a golf ball after it was hit. Write an equation in vertex form that represents the path of the ball. y = –(x – 50)2 + 150 y = –(x – 150)2 + 50 y = –(x – 100)2 + 150 y = –(x – 50)2 + 150

OpenStudy (anonymous):

last one @satellite73

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

my eyes aren't that good, but i looks like the vertex is \((50,150)\) is that right ?

OpenStudy (anonymous):

i cant see it either :/

OpenStudy (anonymous):

if so it is \[y=-(x-50)^2+150\] it looks like you posted that answer twice, first and last

OpenStudy (anonymous):

looks like the top of the parabola is where \(x=50\) and \(y=150\)

OpenStudy (anonymous):

thanks .. Your the bestXD

OpenStudy (anonymous):

go with \[y=-(x-50)^2+150\] first answer and also the last, unless that was a typo

OpenStudy (anonymous):

happy to help. zombies rule

OpenStudy (anonymous):

do you know what your grade was on this?

OpenStudy (anonymous):

@ZombieGirl

OpenStudy (spicytaco1):

oןןǝɥ